The acient Greeks thought that the most pleasing shape for a rectangle was one for which the ratio of the length to the width was 8 to 5, the golden ratio. If the length of the painting is 2 ft longer than its width, then for what dimensions would the length and width have the golden ratio?
Answer by checkley71(8403)

8/5=(X+2)/X NOW CROSS MULTIPLY.
8X=5(X+2)
8X=5X+10
8X-5X=10
3X=10
X=10/3
X=3.333
3.333+2=5.333
THUS THE GOLDEN RATIO HERE IS 3.333 BY 5.333.
PROOF:
8/5=5.333/3.333
1.6=1.6